# Compressible and immiscible fluids with arbitrary density ratios: diffuse interface theory

### v3

@@ I Introduction
  • [II(B)]: Uniform excess volume. According to the definition of density, we have the following expression for the mixture density at any position $\bm{r}$ in the considered domain $\Omega$ as
@@ II.3 Total time derivative, mass conservation derivative, and validity of the continuity equation
We note that, in general, the so-called mass conservative derivative should not be used to derive the kinetic equation. The key issue is that mass conservation has not been connected to the physical principle of dissipation, as will be demonstrated below. Furthermore, the mass conservative derivative is problematic for deriving the continuity equation, as illustrated in the following. In the absence of diffusion, the convective mass flux $\mathbf{j}_{c}$ is commonly expressed in the literature as
@@ III Setup, definition and excess volume concept
The present work develops a generalized theoretical framework for systems with nonzero excess volume within the fluid–fluid interface, particularly for cases involving large density ratios. The objective is to extend both the classical van der Waals and Euler formulations by incorporating interfacial compressibility. Here, compressibility refers to the variation of density across the interface arising from mixing with a finite excess volume. Even when the excess volume is nonzero, if $\rho_{w}^{\prime}\approx\rho_{a}^{\prime}=\rho^{*}$, the single-variable van der Waals formulation remains adequate to represent the compressibility, since $\rho\approx\rho^{*}$ according to Eq. (30). Either the variation of $\rho^{*}$ or $\phi$ can then describe the compressibility. However, when $\rho_{w}^{\prime}\gg\rho_{a}^{\prime}$, the van der Waals formulation loses its validity and must be generalized to properly describe the evolution of the density field.
@@
The boundary of the domain $\Omega$ is denoted by $\Gamma$. For a binary system, we use the notation $\phi_{1}=\phi$, $\phi_{2}=1-\phi$. The density of the mixture at any position $\bm{r}\in\Omega$ is denoted by $\rho(\bm{r})$.
@@ VI.1 Energy formulation and dissipation
The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})\mathrm{d}\Omega.$ (53)
@@
$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
where $\bm{\zeta}$ is the linear momentum density expressed as $\bm{\zeta}=\rho\mathbf{u}$.
@@
Next, we evaluate the total time derivative of the free energy $\mathcal{F}$, the pressure energy $\mathcal{P}$, and the kinetic energy $\mathcal{K}$. By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\bm{\phi}$ and $\nabla\bm{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
@@
$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
@@
$\underline{\underline{\bm{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (62)
The first part of Eq. (60) defines a generalized formulation for the chemical potential, as stated in Eq. (61). At thermodynamic equilibrium, the chemical potential is a constant value throughout the system. In a special case of a zero chemical potential, i.e., $\mu_{i}=0$, Eq. (61) is equivalent to the Euler-Lagrange equation and replicates the Young-Laplace equation; Eq. (61) also provides a physical interpretation for the variational derivative in mathematics. The second part of Eq. (60) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999). This stress tensor is consistent with the Korteweg stress by evaluating the derivative, $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
@@
$\dot{\mathcal{K}}=\int_{\Omega}(\dot{\bm{\zeta}}\cdot\mathbf{u}+\bm{\zeta}\cdot\dot{\mathbf{u}})\mathrm{d}\Omega,$ (64)
@@
$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
@@
$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]\cdot\mathbf{u}+\bigg(\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)
@@ VI.2 Kinetic equations
$\dot{\bm{\zeta}}=\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I}).$ (68)
The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$. This result is nothing but momentum balance. Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\bm{\Theta}}}$. The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy. Note that a viscous dissipation term cannot be added to the momentum conservation equation, as the pressure term does not contain the information of the velocity entropy [see Eq. (58)].
The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\bm{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$. By applying the same dissipation principle as done for the scalar field in Eq. (67), we obtain the dissipation equation for the velocity as
$\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}.$ (69)
Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ [see Eq. (67)], three terms $\nabla\bm{\zeta}$, $\nabla\bm{\zeta}^{T}$, and $(\nabla\cdot\bm{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$. The physical meaning of the velocity dissipation is the increase of the velocity entropy $s_{u}$. Within the framework of entropy, the third term, III, in Eq. (65) can be interpreted as
$\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
@@
$\nabla\cdot\tau_{1}\nabla\bm{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \bm{\zeta}=\partial k/\partial\mathbf{u};$ (72)
@@
The linear momentum, $\bm{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
@@
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\bm{\zeta},$ (74)
@@ VI.3 Density evolution equation and the generalization of Bernoulli’s law
Eq. (68) subtracting Eq. (69) with the relation $\dot{\bm{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\bm{\zeta}+\nabla\bm{\zeta}^{T})+\tau_{2}(\nabla\cdot\bm{\zeta})\mathbf{I}].$ (75)
@@
$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\bm{\Theta}}})].$ (76)
@@ VI.4 The problem of the conventional kinetic energy formulation
A noteworthy shortcoming of Eq. (79) is the missing information of the momentum balance as well as the transformation between the potential and kinetic energies. Moreover, the first part in Eq. (79) requires a prior information about the time evolution of the density, which is unknown for two immiscible phases within the interface, as the excess volume within the interface is unknown. It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution. The reasons are three-fold:
(I) The classical continuity equation is derived for a one-component fluid, in which a single variable-either the density $\rho$ or the volume composition $\phi$-is sufficient to characterize the thermodynamic state. In such systems, density variations with pressure follow a standard equation of state (EOS), such as the van der Waals EOS. However, within the interface between two immiscible fluids, mixing entropy leads to a local excess volume that is generally non-zero and non-uniform in space. Under these conditions, the classical van der Waals EOS loses validity because it applies only to ideal mixing, where the excess volume of two mixing components is zero everywhere. For non-ideal mixing of two components, three volume contributions must be considered: the component volumes $v_{1}$ and $v_{2}$, and the excess volume $v_{e}$. When $v_{e}=0$, the classical EOS may be applied to $v_{1}$ and $v_{2}$, relating the partial pressures $p_{1}$ and $p_{2}$ to the corresponding densities $\rho_{1}$ and $\rho_{2}$. When $v_{e}\neq 0$, an additional EOS governing $v_{e}$ is required, and this cannot be handled within the classical framework of continuity equations because an additional degree of freedom due to excess volume has been introduced. As discussed in Section III, our approach is to redistribute the excess volume $v_{e}$ into $v_{1}$ and $v_{2}$ and to introduce a new volume concentration that incorporates the excess-volume information. With this redistribution, the effective densities $\rho_{1}$ and $\rho_{2}$ are no longer constants; their actual values, including the redistributed excess volume, depend on the chosen redistribution scheme and are not of primary interest here.
Instead, we derive a new density equation that directly relates density variations to changes in both pressure and velocity entropy [Eq. (75)]. This stands in contrast to the classical continuity equation, in which density variations are independent of pressure and entropy. In the classical framework, one must prescribe an EOS externally to obtain the kinetic evolution. In the present theory, by contrast, the EOS emerges naturally from the evolution equations themselves (see Section VIB), aided by the pressure obtained from the modified Gibbs–Duhem relation and the newly defined volume concentration that incorporates the excess volume.
(II) In the case of a spatially uniform density for one-component fluids, the classical continuity equation reduces to $\nabla\cdot\mathbf{u}=0$. This divergence-free condition, which characterizes incompressible one-component fluids, does not hold within the interface between two immiscible fluids. Enforcing $\nabla\cdot\mathbf{u}=0$ in such regions produces a pressure field that lacks physical meaning. When the excess volume is nonzero, the fluid is inherently compressible, and thus deviates from the incompressibility assumption. In this case, the pressure associated with the compressibility arising from the excess volume must be determined from the modified Gibbs–Duhem relation, rather than from solving $\nabla\cdot\mathbf{u}=0$. As already discussed in Section III, two variables are required to describe the thermodynamic state of the fluids. The volume concentration accounts for the compressibility and determines the pressure according to Eq. (58). The density evolution equation then naturally gives rise to EOS.
(III) The classical continuity equation cannot incorporate diffusion processes, such as evaporation and condensation, which are incompatible with the incompressibility assumption, as noted in the Introduction and discussed elsewhere Abels et al. (2012). In fact, the diffusion equation arises from the presence of excess volume or equivalently from the mixing entropy $s_{\phi}$, thereby leading to an effective compressibility.
@@ VI.6 Remarks
In the current work, we consider the linear momentum $\bm{\zeta}=\rho\mathbf{u}$. Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\bm{\zeta}\propto\mathbf{u}^{n}$ or $\bm{\zeta}\propto(\nabla\mathbf{u})^{n}$. The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects. These specialized cases are out of the scope of the present discussion.

### v4

@@ I Introduction
  • [II(B)]: Uniform excess volume. According to the definition of density, we have the following expression for the mixture density at any position $\boldsymbol{r}$ in the considered domain $\Omega$ as
@@ II.3 Total time derivative, mass conservation derivative, and validity of the continuity equation
We note that, in general, the so-called mass conservative derivative should not be used to derive the kinetic equation. The key issue is that mass conservation has not been connected to the physical principle of dissipation, as will be demonstrated below. Furthermore, the mass-conservative derivative poses difficulties for deriving the continuity equation, as illustrated below and discussed in detail in Appendix D. In the absence of diffusion, the convective mass flux $\mathbf{j}_{c}$ is commonly expressed in the literature as
@@ III Setup, definition and excess volume concept
The present work develops a generalized theoretical framework for systems with nonzero excess volume within the fluid–fluid interface, particularly for cases involving large density ratios. The objective is to extend both the classical van der Waals and Euler formulations by incorporating interfacial compressibility. Here, compressibility refers to the variation of density across the interface arising from mixing with a finite excess volume. Even when the excess volume is nonzero, if $\rho_{w}^{\prime}\approx\rho_{a}^{\prime}=\rho^{*}$, the single-variable van der Waals formulation remains adequate to represent the compressibility, since $\rho\approx\rho^{*}$ according to Eq. (30). Either the variation of $\rho^{*}$ or $\phi$ can then describe the compressibility. However, when $\rho_{w}^{\prime}\gg\rho_{a}^{\prime}$, the van der Waals formulation loses validity in the absence of a prescribed equation of state (EOS) and must therefore be generalized to accurately describe the evolution of the density field.
@@
The boundary of the domain $\Omega$ is denoted by $\Gamma$. For a binary system, we use the notation $\phi_{1}=\phi$, $\phi_{2}=1-\phi$. The density of the mixture at any position $\boldsymbol{r}\in\Omega$ is denoted by $\rho(\boldsymbol{r})$.
@@ VI.1 Energy formulation and dissipation
The integration over the local free energy density $e(\boldsymbol{\phi},\nabla\boldsymbol{\phi})$ leads to
$\mathcal{F}(\boldsymbol{\phi},\nabla\boldsymbol{\phi})=\int_{\Omega}e(\boldsymbol{\phi},\nabla\boldsymbol{\phi})\mathrm{d}\Omega.$ (53)
@@
$\mathcal{K}=\int_{\Omega}\boldsymbol{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
where $\boldsymbol{\zeta}$ is the linear momentum density expressed as $\boldsymbol{\zeta}=\rho\mathbf{u}$.
@@
Next, we evaluate the total time derivative of the free energy $\mathcal{F}$, the pressure energy $\mathcal{P}$, and the kinetic energy $\mathcal{K}$. By using the chain rule and noting that the functional $\mathcal{F}$ has two arguments, $\boldsymbol{\phi}$ and $\nabla\boldsymbol{\phi}$, the total time derivative of the potential free energy $\mathcal{F}$ reads
$\dot{\mathcal{F}}(\boldsymbol{\phi},\nabla\boldsymbol{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
@@
$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\boldsymbol{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
@@
$\underline{\underline{\boldsymbol{\Theta}}}=\sum_{i}\frac{\partial e}{\partial\nabla\phi_{i}}\otimes\nabla\phi_{i}.$ (62)
The first part of Eq. (60) defines a generalized formulation for the chemical potential, as stated in Eq. (61). At thermodynamic equilibrium, the chemical potential is a constant value throughout the system. In a special case of a zero chemical potential, i.e., $\mu_{i}=0$, Eq. (61) is equivalent to the Euler-Lagrange equation and replicates the Young-Laplace equation; Eq. (61) also provides a physical interpretation for the variational derivative in mathematics. The second part of Eq. (60) defines a stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999). This stress tensor is consistent with the Korteweg stress by evaluating the derivative, $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
@@
$\dot{\mathcal{K}}=\int_{\Omega}(\dot{\boldsymbol{\zeta}}\cdot\mathbf{u}+\boldsymbol{\zeta}\cdot\dot{\mathbf{u}})\mathrm{d}\Omega,$ (64)
@@
$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I})+\dot{\boldsymbol{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\boldsymbol{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
@@
$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I})+\dot{\boldsymbol{\zeta}}]\cdot\mathbf{u}+\bigg(\boldsymbol{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)
@@ VI.2 Kinetic equations
$\dot{\boldsymbol{\zeta}}=\frac{\mathrm{d}(\rho\mathbf{u})}{\mathrm{d}t}=-\nabla\cdot(\underline{\underline{\boldsymbol{\Theta}}}+p\mathbf{I}).$ (68)
The conservative force at the right hand side is responsible for the exchange of the potential energy $\mathcal{F}+\mathcal{P}$ with the kinetic energy $\mathcal{K}$. This result is nothing but momentum balance. Here, a more clear physical meaning of the force is presented; the force term consists of the pressure $p$ and the stress tensor $\underline{\underline{\boldsymbol{\Theta}}}$. The stress tensor is due to the gradient term in the potential energy $\mathcal{F}$, while the pressure is contributed by the associated free energy. Note that a viscous dissipation term cannot be added to the momentum conservation equation, as the pressure term does not contain the information of the velocity entropy [see Eq. (58)].
The only term left in the time evolution of system free energy, III, is $\int_{\Omega}\boldsymbol{\zeta}\cdot\dot{\mathbf{u}}\mathrm{d}\Omega$. By applying the same dissipation principle as done for the scalar field in Eq. (67), we obtain the dissipation equation for the velocity as
$\dot{\mathbf{u}}=\nabla\cdot\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\nabla\cdot\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}.$ (69)
Unlike the dissipation rule for the scalar variable $\phi_{i}$ with the associated potential $\nabla\mu_{i}$ [see Eq. (67)], three terms $\nabla\boldsymbol{\zeta}$, $\nabla\boldsymbol{\zeta}^{T}$, and $(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}$ have to be taken into account for the dissipation of a vector $\mathbf{u}$. The physical meaning of the velocity dissipation is the increase of the velocity entropy $s_{u}$. Within the framework of entropy, the third term, III, in Eq. (65) can be interpreted as
$\boldsymbol{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}.$ (70)
@@
$\nabla\cdot\tau_{1}\nabla\boldsymbol{\zeta}=\nabla\cdot\tau_{1}\nabla(\partial k/\partial\mathbf{u});\ \boldsymbol{\zeta}=\partial k/\partial\mathbf{u};$ (72)
@@
The linear momentum, $\boldsymbol{\zeta}$, plays a role analogous to that of the chemical potential, $\mu$.
@@
$\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}=-\tau_{0}\boldsymbol{\zeta},$ (74)
@@ VI.3 Density evolution equation and the generalization of Bernoulli’s law
Eq. (68) subtracting Eq. (69) with the relation $\dot{\boldsymbol{\zeta}}=\dot{\rho}\mathbf{u}+\rho\dot{\mathbf{u}}$, we obtain the temporal equation for the density as
$\mathbf{u}\dot{\rho}=-\nabla\cdot[(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})]-\rho\nabla\cdot[\tau_{1}(\nabla\boldsymbol{\zeta}+\nabla\boldsymbol{\zeta}^{T})+\tau_{2}(\nabla\cdot\boldsymbol{\zeta})\mathbf{I}].$ (75)
@@
$\mathbf{u}(\mathbf{u}\cdot\nabla\rho)=-[\nabla\cdot(p\mathbf{I}+\underline{\underline{\boldsymbol{\Theta}}})].$ (76)
@@ VI.4 The problem of the conventional kinetic energy formulation
A noteworthy shortcoming of Eq. (79) is the missing information of the momentum balance as well as the transformation between the potential and kinetic energies. Moreover, the first part in Eq. (79) requires a prior information about the time evolution of the density, which is unknown for two immiscible phases within the interface, as the excess volume within the interface is unknown. It is important to note that the classical continuity equation, $\partial_{t}\rho+\nabla\cdot(\rho\mathbf{u})=0$, cannot be applied here for density evolution (see Appendix D).
@@ VI.6 Remarks
In the current work, we consider the linear momentum $\boldsymbol{\zeta}=\rho\mathbf{u}$. Non-linear terms related to fluid velocity may be accounted for the dissipation if we consider non-linear momentum, for example, $\boldsymbol{\zeta}\propto\mathbf{u}^{n}$ or $\boldsymbol{\zeta}\propto(\nabla\mathbf{u})^{n}$. The non-linear momentum not only modifies the velocity dissipation but also changes the momentum balance, which may be applied for understanding highly nonlinear effects. These specialized cases are out of the scope of the present discussion.